A committee of 9 members is voting on a proposal. Each member casts a yea or nay vote. On a random voting basis, what is the probability that (a) the final vote count is unanimous? (b) the first 4 votes are yea? The last 5 votes are nay?
Answer by edjones(8007)

a)2*0.5^9=.001953*2=.0039
b).5^4=.0625
c).5^5=.0313
.
Ed